A solution with weight percentage of sodium hydroxide 6%is prepared by adding the following mass of water to 209g of that solution with the weight percent of sodium hydroxide of 30% find the amount of water to be added

Grade:8

1 Answers

Aditya
7 years ago
w/W % = 30% (where w is weight of sodium and W is weight of solution)w/209 * 100 = 30w = 30 * 209/100w = 62.7w/(W + x) * 100 = 6 ( where x is mass of water)Substituting we get x = 836 gm